OpenStudy (anonymous):

Will medal and fan 2x-1=4x+3 What is X

OpenStudy (anonymous):

@yomamabf

OpenStudy (anonymous):

@yolomcswagginsggg

OpenStudy (anonymous):

@chosenmatt

OpenStudy (anonymous):

@MikuxLuka

OpenStudy (anonymous):

-2 i think

OpenStudy (yolomcswagginsggg):

Let's solve your equation step-by-step. 2x−1=4x+3 Step 1: Subtract 4x from both sides. 2x−1−4x=4x+3−4x −2x−1=3 Step 2: Add 1 to both sides. −2x−1+1=3+1 −2x=4 Step 3: Divide both sides by -2. −2x −2 = 4 −2 x=−2 Answer: x=−2

OpenStudy (anonymous):

@yolomcswagginsggg thx

OpenStudy (yolomcswagginsggg):

Hey! np